Could you please help me to solve this exercise? If possible, please do it in very details so I can understand where numbers come from, it would help me a lot for other similar exercises

Many many thanks 5. Choosing the relevant information within the following, compute the net trade cycle for the company YEAR 2016 Sales Receivables (ending balance) Inventories (beginning balance) Inventories (ending balance) Accounts payables (ending balance) Cost of goods sold Purchases Interest rate Net financial obligations 800.000,00 110.000,00 120.000,00 64.000,00 840.000,00 600.000,00 10% 400.000,00 Which is the relation between the working capital requirement and the net trade cycle?
Answer #1

Net trade cycle=days sales outstanding+days inventory outstanding-days account payable
days sales outstanding=(365* account receivable)/sales
=(365*110000)/800000
=50.19 days
days inventory outstanding=(365*avg inventory)/cost of goods sold
=(365*((220000+120000)/2))/840000
=73.87 days
days account payable=(365* account payable)/cost of goods sold
=(365*64000)/840000
=27.81 days
Net trade cycle=50.19+73.87-27.81=96.25 days
working capital and net trade cycle are not directky related for well being of company both are important